OpenStudy (anonymous):

The table below shows the ticket rates for whale watching trips offered by Ocean Tours: Age Ticket Price Under 3 years free 3 to 12 years $36 Over 12 years $48 On a certain day, the company took 92 people on whale watching trips. There were 40 children aged 12 and under, of which some children were under 3 years. If x represents the number of children under 3 years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day?

OpenStudy (anonymous):

Here are the answers/choices: (92 - x)36 + (92 - 40) ⋅ 48 = C (40 + x)36 + (92 - (40 + x)) ⋅ 48 = C (92 + x)36 + (92 - (40 + x)) ⋅ 48 = C (40 - x)36 + (92 - 40) ⋅ 48 = C

OpenStudy (yttrium):

I believe it is D.

OpenStudy (anonymous):

And your explanation?

OpenStudy (anonymous):

Thanks, I'll go with that I hope it's right . . .

OpenStudy (yttrium):

(40-x)36 represents the total amount collected from 3-12 yrs old since x represents the no of children under 3. (92-40)48 are the money collected from children over 12 yrs since there are only 40 children 12 and under. Hope you get it. :))

OpenStudy (anonymous):

Ah, I get it, sounds like you're right.

OpenStudy (anonymous):

It's a 22 question test and I'm only on Question 5. So I won't find out anytime soon xD

OpenStudy (yttrium):

you can post it, now. And maybe I can still help but may be I should log-out later.

OpenStudy (anonymous):

Post what? The entire test? Well I don't need help on every single question but for the harder ones I don't know I'll ask you, okay?